# PlainMamba: Improving Non-Hierarchical Mamba in Visual Recognition

### v1

@@ 4.2 COCO object detection and instance segmentation
Setting: Following [100], we test PlainMamba’s ability on COCO object detection and instance segmentation using both the two-stage Mask R-CNN [38] and the single-stage RetinaNet [53]. For both models, we report the results of both 1$\times$ schedule. Following [100], we use ViTAdapter [11] to compute multi-scale features to fit the FPN network structure. We use the commonly used training settings proposed in [59] to keep a fair comparison.
@@
Backbone Hierarchical Params FLOPs $AP^{bb}$ $AP^{bb}_{50}$ $AP^{bb}_{75}$ CNN ResNeXt101-32x4d [96] 56M 319G 39.9 - - ResNeXt101-64×4d [96] 95M 413G 41.0 - - Transformer Swin-Tiny [59] 38M 245G 41.5 - - Swin-Small [59] 60M 335G 44.5 - - Focal-Tiny [101] 39M 265G 43.7 - - Focal-Small [101] 62M 367G 45.6 - - PVT-Small [86] 34M - 40.4 61.3 43.0 PVT-Medium [86] 54M - 41.9 63.1 44.3 PVT-Large [86] 71M - 42.6 63.7 45.4 State Space Modeling EfficientVMamba-T [64] 13M - 37.5 57.8 39.6 EfficientVMamba-S [64] 19M - 39.1 60.3 41.2 EfficientVMamba-B [64] 44M - 42.8 63.9 45.8 PlainMamba-Adapter-L1 19M 250G 41.7 62.1 44.4 PlainMamba-Adapter-L2 40M 392G 43.9 64.9 47.0 PlainMamba-Adapter-L3 67M 478G 44.8 66.0 47.9
Results: We report the results of Mask R-CNN object detection and instance segmentation in Table 3 and RetinaNet object detection in Table 4. For Mask R-CNN, PlainMamba is on par with the state-of-the-art architectures. For example, with similar FLOPs and many fewer parameters, PlainMamba-L1 achieves 44.1 APbb and 39.1 APmk when using 1$\times$ training schedule, while Swin-Small achieves 44.8 APbb and 40.9 APmk. We also observe that hierarchical models tend to work better than non-hierarchical models. Although our model achieves lower performance than some hierarchical models, e.g., the concurrent VMamba [58], PlainMamba achieves the best performance among its non-hierarchical counterparts. For instance, when using 1$\times$ training schedule, PlainMamba achieves 3.1 higher APbb and 1.6 higher APmk than DeiT-T when they are both equipped with the ViTAdapter [11]. These results demonstrate that PlainMamba is able to extract good local features, which is important to the object-level tasks like instance segmentation. Similarly, PlainMamba also performs well with the single-stage RetinaNet object detector. For example, with only half the model size and similar FLOPs, PlainMamba-L1 achieves 0.2 higher AP than Swin-Tiny. On the other hand, we also admit that PlainMamba is performing worse than the hierarchical VMamba [58]. We attribute such inferiority to the multi-resolution architecture of FPN-based [52] Mask R-CNN, which is more naturally suitable to the hierarchical designs.
Table 5: ADE20K semantic segmentation using UperNet. The FLOPs are computed using input size 512$\times$2048. PlainMamba has much smaller FLOPs here because we used feature maps in a constant resolution of 16.
Backbone Hierarchical Params FLOPs mIoU CNN ResNet-50 [39] 67M 953G 42.1 ResNet-101 [39] 85M 1030G 44.0 ConvNeXt-T [60] 60M 939G 46.7 Transformer DeiT-S+MLN [80] 58M 1217G 43.8 DeiT-B+MLN [80] 144M 2007G 45.5 XCiT-T12/8 [2] 34M - 43.5 XCiT-S12/8 [2] 52M 1237G 46.6 XCiT-S24/8 [2] 74M 1587G 48.1 Swin-Tiny [59] 60M 945G 44.5 Swin-Small [59] 81M 1038G 47.6 Focal-Tiny [101] 62M 998G 45.8 Focal-Small [101] 85M 1130G 48.0 Twins-SVT-Small [14] 54M 912G 46.2 Twins-SVT-Small [14] 88M 1044G 47.7 State Space Modeling ViM-T [108] 13M - 41.0 ViM-S [108] 46M - 44.9 LocalVim-T [45] 36M 181G 43.4 LocalVim-S [45] 58M 297G 46.4 VMamba-T [58] 55M 964G 47.3 VMamba-S [58] 76M 1081G 49.5 PlainMamba-L1 35M 174G 44.1 PlainMamba-L2 55M 285G 46.8 PlainMamba-L3 81M 419G 49.1
4.3 ADE20K Semantic Segmentation
Setting: We follow common practice [59, 23, 100] and benchmark PlainMamba’s ability on semantic segmentation with UperNet [94] as the segmentation network. Unlike XCiT [2], we do not explicitly resize the constant resolution feature maps into multi-scale. Following [59], we train all models for 160 iterations with batch size 16 and set the default training image size to 512$\times$512.
Results: We report our model’s ADE20K semantic segmentation performance in Table 5. Similar to the ImageNet-1k and COCO experiments, here the competing models include both hierarchical and non-hierarchical backbones in three types of visual feature extractors. The results again suggest that PlainMamba achieves the best performance among the non-hierarchical models. For example, with similar parameter amounts, PlainMamba-L2 outperforms the high-resolution (patch size of 8) XCiT-S12/8 model [2] with a much lower computation cost. Moreover, PlainMamba-L2 also outperforms the hierarchical Swin-Transformer-Tiny [59], achieving better mIoU while having a lower model size and FLOPs. At the same time, PlainMamba is also doing better than the concurrent Vision Mamba [108]. For instance, PlainMamba-L2 achieves a 1.9 higher mIoU than ViM-S. This result verifies our model’s effectiveness in extracting fine-grained visual features, which is essential for the pixel-wise semantic segmentation task.
4.4 Ablation Studies and Discussions
@@
Setting: Here, we conduct ablation studies to test our model designs and to gain a deeper understanding of the proposed method. We use our L1 model, with less than 10M parameters, for most experiments. The models are all pre-trained on ImageNet-1K following the same training settings described in Section 4.1.
Depth v.s. Width When designing neural architectures for a given parameter count, it’s usually important to find a good balance between the network’s depth, i.e., the number of layers, and its width, i.e., the feature dimensions. While this problem was studied for existing architectures [93, 99], it is still unclear whether the previous conclusions are applicable to vision SSMs. In Table 6, we study the depth and width trade-off of the proposed PlainMamba. Firstly, the results show that deeper models tend to perform better than shallow ones. For example, when the parameter count is around 7.4M, the 12-layer model achieves 2.2% higher ImageNet top-1 accuracy than the 6-layer counterparts, and the 24-layer model is further 1.1% higher than the 12-layer one. However, when we further increase the depth to 36 while reducing the width accordingly, the top-1 accuracy remains similar. On the other hand, we also notice that deeper models are less efficient than shallower but wider models. For instance, the 24-layer model is 0.3G FLOPs higher than the 12-layer model. These results suggest the necessity of a good balance between network depth and width.
Table 7: Ablation study of model depth v.s. width on ImageNet-1K.
Method Params FLOPs Top-1 VisionMamba [108] 7.8M 1.3G 74.4 VMamba [58] 7.3M 3.0G 77.1 Ours 7.3M 3.0G 77.9
PlainMamba Block Design. Here, we test different designs of the PlainMamba block by comparing it with the block designs in Vision Mamba [108] and VMamba [58]. For a fair comparison, we use the same model depth and width settings for all designs and train all models with the same training recipe. We also remove the CLS tokens from the Vision Mamba [108] block and use the global averaging pooling as an alternative. We report the results in Table 7. We can see that our design achieves the best results. Specifically, Vision Mamba only achieves a 74.4% ImageNet accuracy, which is 3.5% lower than ours. We also notice that the model with a Vision Mamba block is inferior to the original Vision Mamba model, which is caused by the removal of the CLS token. These results suggest that our model still retains its ability when the CLS token is absent. Also, our design performs better than the VMamba block [58] with a 0.8% accuracy advantage, indicating that the improvements come from our proposed Continuous 2D Scanning and Direction-aware Updating, which validate the effectiveness of our proposed techniques in adapting SSM for 2D images.
@@
Efficiency Comparison with ViT One particular advantage of SSMs, e.g., Mamba, is their ability to capture global information while maintaining efficiency. in Figure 4, we compare the PlainMamba’s efficiency with the vision transformer. Specifically, to ensure a fair comparison, we create a DeiT model with channel numbers of 224, resulting in a model with 7.4M parameters, which is used to compare with PlainMamba-L1. Specifically, we compare the model FLOPs and the peak inference memory using inputs of different sizes. The results show that our model is able to keep the computation cost low when the input size is scaled up to high resolutions, e.g., 4096$\times$4096. However, DeiT’s FLOPs and memory consumption increase rapidly when using such high-resolution inputs. On the other hand, we also notice that our model’s efficiency is inferior to the similar-sized DeiT when using low-resolution images, e.g., 128$\times$128. To further investigate such a difference in their efficiency, we decompose their FLOPs into three parts [79]: 1) token mixing, 2) channel mixing, and 3) others. Specifically, token mixing refers to the multi-head attention part in DeiT and the selective scanning part in PlainMamba, and channel mixing refers to the feed-forward network in DeiT and the input & output projection in PlainMamba. We report the results in Table 8. These suggest that PlainMamba’s FLOPs are evenly distributed across the three parts in low and high resolutions. On the contrary, when using 128$\times$128 inputs, DeiT’s FLOPs are dominated by channel-mixing and the other two parts are negligible. However, because of the quadratic complexity of self-attention operation, DeiT’s FLOPs in token mixing grow to 23T when using 4096$\times$4096 inputs, 23 times more expensive than PlainMamba. These results verify PlainMamba’s high efficiency when using high-resolution inputs.
5 Conclusion
In this paper, we present PlainMamba, a plain SSM-based model for visual recognition. Our model is conceptually simple because it does not use any special tokens and does not have a hierarchical structure, making it a perfect counterpart to the widely used plain vision transformer. To better adapt SSM to model 2D visual features, we propose two techniques, namely (i) Continuous 2D Scanning and (ii) Direction-aware updating. We benchmark PlainMamba on a wide range of visual recognition tasks, including image classification, object detection, instance segmentation, and semantic segmentation. The results show that PlainMamba achieves superior performance to previous non-hierarchical models, including the concurrent SSM-based models, and can perform on par with the high-performing hierarchical models. Moreover, when being used for high-resolution inputs, PlainMamba requires significantly less computing compared to the vision transformer. We hope our model can serve as a baseline method for future research in this area.
Acknowledgements

### v2

@@ 4.1 Experiment Settings
COCO Object Detection and Instance Segmentation. Following [Yang et al.(2023)Yang, Xu, Mello, Crowley, and Wang], we test PlainMamba’s ability on COCO object detection and instance segmentation using both the two-stage Mask R-CNN [He et al.(2017)He, Gkioxari, Dollar, and Girshick] and the single-stage RetinaNet [Lin et al.(2017b)Lin, Goyal, Girshick, He, and Dollar]. For both models, we report the results of both 1$\times$ schedule. Following [Yang et al.(2023)Yang, Xu, Mello, Crowley, and Wang], we use ViTAdapter [Chen et al.(2022b)Chen, Duan, Wang, He, Lu, Dai, and Qiao] to compute multi-scale features to fit the FPN network structure. We use the commonly used training settings proposed in [Liu et al.(2021)Liu, Lin, Cao, Hu, Wei, Zhang, Lin, and Guo] to keep a fair comparison.
4.2 Main Results
ImageNet-1K Classification. In Table 2, we report the ImageNet-1K experiment results. We compare PlainMamba with three different kinds of visual feature extractors: CNNs, vision transformers, and SSMs. In addition, the comparison includes both hierarchical and non-hierarchical models. Firstly, when comparing with SSMs, our model is doing better than the recently proposed Vision Mamba [Zhu et al.(2024)Zhu, Liao, Zhang, Wang, Liu, and Wang] and Mamba-ND [Li et al.(2024)Li, Singh, and Grover]. For example, PlainMamba-L2 achieves a 2.4% higher accuracy than Mamba-ND-T while they share a similar model size. These results validate PlainMamba’s effectiveness as a non-hierarchical SSM. Secondly, when compared with CNNs and transformers, our model achieves better performance than the non-hierarchical counterparts. For example, PlainMamba-L2 achieves 1.7% better accuracy with DeiT-Small. Moreover, PlainMamba also achieves similar performance when compared with hierarchical models. For example, when the model size is around 25M, our model achieves 0.3% better accuracy than Swin-Tiny, validating PlainMamba’s ability as a general feature extractor. On the other hand, the hierarchical VMamba [Liu et al.(2024c)Liu, Tian, Zhao, Yu, Xie, Wang, Ye, and Liu], together with other hierarchical transformers, do achieve a better accuracy than ours. As we explained in Section 1, hierarchical models tend to perform better than non-hierarchical ones in visual recognition. As the main motivation of our work is to develop a simple Mamba architecture, a bit inferior ImageNet accuracy is acceptable.
Table 3: ADE20K semantic segmentation using UperNet. The FLOPs are computed using input size 512$\times$2048.
Backbone Hierarchical Params FLOPs mIoU CNN ResNet-50 [He et al.(2016)He, Zhang, Ren, and Sun] 67M 953G 42.1 ResNet-101 [He et al.(2016)He, Zhang, Ren, and Sun] 85M 1030G 44.0 ConvNeXt-T [Liu et al.(2022)Liu, Mao, Wu, Feichtenhofer, Darrell, and Xie] 60M 939G 46.7 Transformer DeiT-S+MLN [Touvron et al.(2021)Touvron, Cord, Douze, Massa, Sablayrolles, and Jégou] 58M 1217G 43.8 DeiT-B+MLN [Touvron et al.(2021)Touvron, Cord, Douze, Massa, Sablayrolles, and Jégou] 144M 2007G 45.5 XCiT-T12/8 [Ali et al.(2021)Ali, Touvron, Caron, Bojanowski, Douze, Joulin, Laptev, Neverova, Synnaeve, Verbeek, et al.] 34M - 43.5 XCiT-S12/8 [Ali et al.(2021)Ali, Touvron, Caron, Bojanowski, Douze, Joulin, Laptev, Neverova, Synnaeve, Verbeek, et al.] 52M 1237G 46.6 XCiT-S24/8 [Ali et al.(2021)Ali, Touvron, Caron, Bojanowski, Douze, Joulin, Laptev, Neverova, Synnaeve, Verbeek, et al.] 74M 1587G 48.1 Swin-Tiny [Liu et al.(2021)Liu, Lin, Cao, Hu, Wei, Zhang, Lin, and Guo] 60M 945G 44.5 Swin-Small [Liu et al.(2021)Liu, Lin, Cao, Hu, Wei, Zhang, Lin, and Guo] 81M 1038G 47.6 Focal-Tiny [Yang et al.(2021)Yang, Li, Zhang, Dai, Xiao, Yuan, and Gao] 62M 998G 45.8 Focal-Small [Yang et al.(2021)Yang, Li, Zhang, Dai, Xiao, Yuan, and Gao] 85M 1130G 48.0 Twins-SVT-Small [Chu et al.(2021)Chu, Tian, Wang, Zhang, Ren, Wei, Xia, and Shen] 54M 912G 46.2 Twins-SVT-Small [Chu et al.(2021)Chu, Tian, Wang, Zhang, Ren, Wei, Xia, and Shen] 88M 1044G 47.7 State Space Modeling ViM-T [Zhu et al.(2024)Zhu, Liao, Zhang, Wang, Liu, and Wang] 13M - 41.0 ViM-S [Zhu et al.(2024)Zhu, Liao, Zhang, Wang, Liu, and Wang] 46M - 44.9 LocalVim-T [Huang et al.(2024)Huang, Pei, You, Wang, Qian, and Xu] 36M 181G 43.4 LocalVim-S [Huang et al.(2024)Huang, Pei, You, Wang, Qian, and Xu] 58M 297G 46.4 VMamba-T [Liu et al.(2024c)Liu, Tian, Zhao, Yu, Xie, Wang, Ye, and Liu] 55M 964G 47.3 VMamba-S [Liu et al.(2024c)Liu, Tian, Zhao, Yu, Xie, Wang, Ye, and Liu] 76M 1081G 49.5 PlainMamba-L1 35M 174G 44.1 PlainMamba-L2 55M 285G 46.8 PlainMamba-L3 81M 419G 49.1
ADE20K Semantic Segmentation We report our model’s ADE20K semantic segmentation performance in Table 3. Similar to the ImageNet-1k and COCO experiments, here the competing models include both hierarchical and non-hierarchical backbones in three types of visual feature extractors. The results again suggest that PlainMamba achieves the best performance among the non-hierarchical models. For example, with similar parameter amounts, PlainMamba-L2 outperforms the high-resolution (patch size of 8) XCiT-S12/8 model [Ali et al.(2021)Ali, Touvron, Caron, Bojanowski, Douze, Joulin, Laptev, Neverova, Synnaeve, Verbeek, et al.] with a much lower computation cost. Moreover, PlainMamba-L2 also outperforms the hierarchical Swin-Transformer-Tiny [Liu et al.(2021)Liu, Lin, Cao, Hu, Wei, Zhang, Lin, and Guo], achieving better mIoU while having a lower model size and FLOPs. At the same time, PlainMamba is also doing better than the concurrent Vision Mamba [Zhu et al.(2024)Zhu, Liao, Zhang, Wang, Liu, and Wang]. For instance, PlainMamba-L2 achieves a 1.9 higher mIoU than ViM-S. This result verifies our model’s effectiveness in extracting fine-grained visual features, which is essential for the pixel-wise semantic segmentation task.
Table 4: Mask R-CNN object detection and instance segmentation on MS COCO mini-val using 1$\times$ schedule. We use ViTAdapter [Chen et al.(2022b)Chen, Duan, Wang, He, Lu, Dai, and Qiao] to compute multi-scale features. FLOPs are computed using input size 1280$\times$800.
Backbone Hierarchical Params FLOPs $AP^{bb}$ $AP^{bb}_{50}$ $AP^{bb}_{75}$ $AP^{mk}$ $AP^{mk}_{50}$ $AP^{mk}_{75}$ CNN ResNeXt101-32x4d [Xie et al.(2017)Xie, Girshick, Dollár, Tu, and He] 63M 340G 41.9 - - 37.5 - - ResNeXt101-64x4d [Xie et al.(2017)Xie, Girshick, Dollár, Tu, and He] 102M 493G 42.8 - - 38.4 - - Transformer ViT-Adapter-T [Chen et al.(2022b)Chen, Duan, Wang, He, Lu, Dai, and Qiao] 29M 349G 41.1 62.5 44.3 37.5 59.7 39.9 ViT-Adapter-S [Chen et al.(2022b)Chen, Duan, Wang, He, Lu, Dai, and Qiao] 49M 463G 44.7 65.8 48.3 39.9 62.5 42.8 ViT-Adapter-B [Chen et al.(2022b)Chen, Duan, Wang, He, Lu, Dai, and Qiao] 131M 838G 47.0 68.2 51.4 41.8 65.1 44.9 PVT-Small [Wang et al.(2021)Wang, Xie, Li, Fan, Song, Liang, Lu, Luo, and Shao] 44M - 40.4 62.9 43.8 37.8 60.1 40.3 PVT-Medium [Wang et al.(2021)Wang, Xie, Li, Fan, Song, Liang, Lu, Luo, and Shao] 64M - 42.0 64.4 45.6 39.0 61.6 42.1 PVT-Large [Wang et al.(2021)Wang, Xie, Li, Fan, Song, Liang, Lu, Luo, and Shao] 81M - 42.9 65.0 46.6 39.5 61.9 42.5 Swin-Tiny [Liu et al.(2021)Liu, Lin, Cao, Hu, Wei, Zhang, Lin, and Guo] 48M 264G 42.2 - - 39.1 - - Swin-Small [Liu et al.(2021)Liu, Lin, Cao, Hu, Wei, Zhang, Lin, and Guo] 69M 354G 44.8 - - 40.9 - - ViL-Tiny [Zhang et al.(2021)Zhang, Dai, Yang, Xiao, Yuan, Zhang, and Gao] 26M 145G 41.4 63.5 45.0 38.1 60.3 40.8 ViL-Small [Zhang et al.(2021)Zhang, Dai, Yang, Xiao, Yuan, Zhang, and Gao] 45M 218G 44.9 67.1 49.3 41.0 64.2 44.1 ViL-Medium [Zhang et al.(2021)Zhang, Dai, Yang, Xiao, Yuan, Zhang, and Gao] 60M 293G 47.6 69.8 52.1 43.0 66.9 46.6 State Space Modeling EfficientVMamba-T [Pei et al.(2024)Pei, Huang, and Xu] 11M 60G 35.6 57.7 38.0 33.2 54.4 35.1 EfficientVMamba-S [Pei et al.(2024)Pei, Huang, and Xu] 31M 197G 39.3 61.8 42.6 36.7 58.9 39.2 EfficientVMamba-B [Pei et al.(2024)Pei, Huang, and Xu] 53M 252G 43.7 66.2 47.9 40.2 63.3 42.9 VMamba-T [Liu et al.(2024c)Liu, Tian, Zhao, Yu, Xie, Wang, Ye, and Liu] 42M 262G 46.5 68.5 50.7 42.1 65.5 45.3 VMamba-S [Liu et al.(2024c)Liu, Tian, Zhao, Yu, Xie, Wang, Ye, and Liu] 64M 357G 48.2 69.7 52.5 43.0 66.6 46.4 PlainMamba-Adapter-L1 31M 388G 44.1 64.8 47.9 39.1 61.6 41.9 PlainMamba-Adapter-L2 53M 542G 46.0 66.9 50.1 40.6 63.8 43.6 PlainMamba-Adapter-L3 79M 696G 46.8 68.0 51.1 41.2 64.7 43.9
COCO Object Detection and Instance Segmentation. We report the results of Mask R-CNN object detection and instance segmentation in Table 4. With similar FLOPs and many fewer parameters, PlainMamba-L1 achieves 44.1 APbb and 39.1 APmk when using 1$\times$ training schedule, while Swin-Small achieves 44.8 APbb and 40.9 APmk. We also observe that hierarchical models tend to work better than non-hierarchical models. Although our model achieves lower performance than some hierarchical models, e.g., the concurrent VMamba [Liu et al.(2024c)Liu, Tian, Zhao, Yu, Xie, Wang, Ye, and Liu], PlainMamba achieves the best performance among its non-hierarchical counterparts. For instance, when using 1$\times$ training schedule, PlainMamba achieves 3.1 higher APbb and 1.6 higher APmk than DeiT-T when they are both equipped with the ViTAdapter [Chen et al.(2022b)Chen, Duan, Wang, He, Lu, Dai, and Qiao]. These results demonstrate that PlainMamba is able to extract good local features, which is important to the object-level tasks like instance segmentation. On the other hand, we also admit that PlainMamba is performing worse than the hierarchical VMamba [Liu et al.(2024c)Liu, Tian, Zhao, Yu, Xie, Wang, Ye, and Liu]. We attribute such inferiority to the multi-resolution architecture of FPN-based [Lin et al.(2017a)Lin, Dollár, Girshick, He, Hariharan, and Belongie] Mask R-CNN, which is more naturally suitable to the hierarchical designs.
5 Conclusion
We present PlainMamba, a plain SSM-based model for visual recognition. Our model is conceptually simple because it uses no special tokens or hierarchical structure, making it a perfect counterpart to the widely used plain vision transformer. The results show that PlainMamba achieves superior performance to previous non-hierarchical models, including the concurrent SSM-based models, and can perform on par with the high-performing hierarchical models. We hope our model can serve as a baseline for future research in this area.
Acknowledgements
Appendix A More Experiments
A.1 COCO Object Detection using RetinaNet
We report COCO RetinaNet object detection in Table 5. Similar to Mask R-CNN, whose results are reported in the main paper, PlainMamba also performs well with the single-stage RetinaNet object detector. For example, with only half the model size and similar FLOPs, PlainMamba-L1 achieves 0.2 higher AP than Swin-Tiny.
@@ A.1 COCO Object Detection using RetinaNet
Backbone Hierarchical Params FLOPs $AP^{bb}$ $AP^{bb}_{50}$ $AP^{bb}_{75}$ CNN ResNeXt101-32x4d [Xie et al.(2017)Xie, Girshick, Dollár, Tu, and He] 56M 319G 39.9 - - ResNeXt101-64×4d [Xie et al.(2017)Xie, Girshick, Dollár, Tu, and He] 95M 413G 41.0 - - Transformer Swin-Tiny [Liu et al.(2021)Liu, Lin, Cao, Hu, Wei, Zhang, Lin, and Guo] 38M 245G 41.5 - - Swin-Small [Liu et al.(2021)Liu, Lin, Cao, Hu, Wei, Zhang, Lin, and Guo] 60M 335G 44.5 - - Focal-Tiny [Yang et al.(2021)Yang, Li, Zhang, Dai, Xiao, Yuan, and Gao] 39M 265G 43.7 - - Focal-Small [Yang et al.(2021)Yang, Li, Zhang, Dai, Xiao, Yuan, and Gao] 62M 367G 45.6 - - PVT-Small [Wang et al.(2021)Wang, Xie, Li, Fan, Song, Liang, Lu, Luo, and Shao] 34M - 40.4 61.3 43.0 PVT-Medium [Wang et al.(2021)Wang, Xie, Li, Fan, Song, Liang, Lu, Luo, and Shao] 54M - 41.9 63.1 44.3 PVT-Large [Wang et al.(2021)Wang, Xie, Li, Fan, Song, Liang, Lu, Luo, and Shao] 71M - 42.6 63.7 45.4 State Space Modeling EfficientVMamba-T [Pei et al.(2024)Pei, Huang, and Xu] 13M - 37.5 57.8 39.6 EfficientVMamba-S [Pei et al.(2024)Pei, Huang, and Xu] 19M - 39.1 60.3 41.2 EfficientVMamba-B [Pei et al.(2024)Pei, Huang, and Xu] 44M - 42.8 63.9 45.8 PlainMamba-Adapter-L1 19M 250G 41.7 62.1 44.4 PlainMamba-Adapter-L2 40M 392G 43.9 64.9 47.0 PlainMamba-Adapter-L3 67M 478G 44.8 66.0 47.9
A.2 Ablation Studies and Discussions
Setting: Here, we conduct ablation studies to test our model designs and to gain a deeper understanding of the proposed method. We use our L1 model, with less than 10M parameters, for most experiments. The models are all pre-trained on ImageNet-1K following the same training settings described in the main paper.
Table 6: Ablation study of model depth v.s. width on ImageNet-1K..
Depth v.s. Width When designing neural architectures for a given parameter count, it’s usually important to find a good balance between the network’s depth, i.e., the number of layers, and its width, i.e., the feature dimensions. While this problem was studied for existing architectures [Wu et al.(2019)Wu, Shen, and Van Den Hengel, Xue et al.(2022)Xue, Shi, Wei, Lou, Liu, and You], it is still unclear whether the previous conclusions are applicable to vision SSMs. In Table 6, we study the depth and width trade-off of the proposed PlainMamba. Firstly, the results show that deeper models tend to perform better than shallow ones. For example, when the parameter count is around 7.4M, the 12-layer model achieves 2.2% higher ImageNet top-1 accuracy than the 6-layer counterparts, and the 24-layer model is further 1.1% higher than the 12-layer one. However, when we further increase the depth to 36 while reducing the width accordingly, the top-1 accuracy remains similar. On the other hand, we also notice that deeper models are less efficient than shallower but wider models. For instance, the 24-layer model is 0.3G FLOPs higher than the 12-layer model. These results suggest the necessity of a good balance between network depth and width.
@@ A.2 Ablation Studies and Discussions
Method Params FLOPs Top-1 VisionMamba [Zhu et al.(2024)Zhu, Liao, Zhang, Wang, Liu, and Wang] 7.8M 1.3G 74.4 VMamba [Liu et al.(2024c)Liu, Tian, Zhao, Yu, Xie, Wang, Ye, and Liu] 7.3M 3.0G 77.1 Ours 7.3M 3.0G 77.9
PlainMamba Block Design. Here, we test different designs of the PlainMamba block by comparing it with the block designs in Vision Mamba [Zhu et al.(2024)Zhu, Liao, Zhang, Wang, Liu, and Wang] and VMamba [Liu et al.(2024c)Liu, Tian, Zhao, Yu, Xie, Wang, Ye, and Liu]. For a fair comparison, we use the same model depth and width settings for all designs and train all models with the same training recipe. We also remove the CLS tokens from the Vision Mamba [Zhu et al.(2024)Zhu, Liao, Zhang, Wang, Liu, and Wang] block and use the global averaging pooling as an alternative. We report the results in Table 7. We can see that our design achieves the best results. Specifically, Vision Mamba only achieves a 74.4% ImageNet accuracy, which is 3.5% lower than ours. We also notice that the model with a Vision Mamba block is inferior to the original Vision Mamba model, which is caused by the removal of the CLS token. These results suggest that our model still retains its ability when the CLS token is absent. Also, our design performs better than the VMamba block [Liu et al.(2024c)Liu, Tian, Zhao, Yu, Xie, Wang, Ye, and Liu] with a 0.8% accuracy advantage, indicating that the improvements come from our proposed Continuous 2D Scanning and Direction-aware Updating, which validate the effectiveness of our proposed techniques in adapting SSM for 2D images.
@@
Efficiency Comparison with ViT One particular advantage of SSMs, e.g., Mamba, is their ability to capture global information while maintaining efficiency. in Figure 4, we compare the PlainMamba’s efficiency with the vision transformer. Specifically, to ensure a fair comparison, we create a DeiT model with channel numbers of 224, resulting in a model with 7.4M parameters, which is used to compare with PlainMamba-L1. Specifically, we compare the model FLOPs and the peak inference memory using inputs of different sizes. The results show that our model is able to keep the computation cost low when the input size is scaled up to high resolutions, e.g., 4096$\times$4096. However, DeiT’s FLOPs and memory consumption increase rapidly when using such high-resolution inputs. On the other hand, we also notice that our model’s efficiency is inferior to the similar-sized DeiT when using low-resolution images, e.g., 128$\times$128. To further investigate such a difference in their efficiency, we decompose their FLOPs into three parts [Tolstikhin et al.(2021)Tolstikhin, Houlsby, Kolesnikov, Beyer, Zhai, Unterthiner, Yung, Steiner, Keysers, Uszkoreit, et al.]: 1) token mixing, 2) channel mixing, and 3) others. Specifically, token mixing refers to the multi-head attention part in DeiT and the selective scanning part in PlainMamba, and channel mixing refers to the feed-forward network in DeiT and the input & output projection in PlainMamba. We report the results in Table 8. These suggest that PlainMamba’s FLOPs are evenly distributed across the three parts in low and high resolutions. On the contrary, when using 128$\times$128 inputs, DeiT’s FLOPs are dominated by channel-mixing and the other two parts are negligible. However, because of the quadratic complexity of self-attention operation, DeiT’s FLOPs in token mixing grow to 23T when using 4096$\times$4096 inputs, 23 times more expensive than PlainMamba. These results verify PlainMamba’s high efficiency for high-resolution inputs.